# Practical Secure Delegated Linear Algebra with Trapdoored Matrices

### v2

@@ 7.2 Security and Correctness Guarantees
Our protocol makes three guarantees. First, in the semi-honest model, the protocol will return exactly $AB$ with probability 1. Second, the protocol guarantees computational zero-knowledge against even a dishonest server. Third, it guarantees a weaker form of correctness: that a server engaged in dishonesty in any setup query or fraction $\alpha$ of the online invocations can be detected with high probability with $O\left(\frac{1}{\alpha}\right)$ additional queries (and if $l$ and $|R|$ are large, much more efficient per-call checks are available via Mat-Mul-Check).
The first guarantee follows from expanding the algebra in the return line and noting that all products except $AB$ cancel. We emphasize that all LPN noise exactly cancels; unlike some LWE-based protocols, we introduce no error in the output.
For the security guarantees, we first notice that Mat-Mul-Check, a procedure first defined (to the best of our knowledge) in [Moh11], will return true if $AB=C$, but will return true with probability $|R|^{-\lambda^{\prime}}$ otherwise. Note that this provides low soundness error at very low computational cost for large rings (assuming that $b=O(a,c)$): for 32-bit integers, for instance, we achieve soundness error $2^{-128}$ when $\lambda^{\prime}=4$. Assuming that $\lambda^{\prime}$ is picked so that this probability is negligible, we demonstrate that Algorithm 2 is CZK.
@@
Given the LPN assumption of Section 5, Algorithm 2 is computational zero-knowledge for the server.
@@
We construct a server-view simulator. The simulator is quite simple: it runs a copy of the client with inputs $A=0,B_{i}=0$. The $L_{i}$ generated by the simulator are identically distributed to those generated by the true client, as those are random. If the server reports any of the products $L_{1}L_{2}\dots L_{i}$ or $ML_{1}L_{2}\dots L_{i}$ falsely, both the simulator and a true client will detect the error and terminate the protocol with probability $\geq 1-|R|^{\lambda^{\prime}}$, so the complete views in this case are statistically indistinguishable. Otherwise, the rest of the client’s messages to the server consist exclusively of matrices taking the form of a hidden matrix $H$ plus an independently sampled matrix of the form $(L_{1}L_{2}\dots L_{d})R+\sum_{i=1}^{d}(L_{1}L_{2}\dots L_{i-1})S_{i}$ (where $R$ and $S_{i}$ are sampled as in Right-Hidden-Mult). The simulated client’s messages are identically composed, except that all of the matrices $H$ are zero. In other words, the total server view contains no information not fixed by $L_{1},\dots,L_{n}$ and a set of matrices decomposable into vectors generated as $h+(L_{1}L_{2}\dots L_{d})r+\sum_{i=1}^{d}(L_{1}L_{2}\dots L_{i-1})s_{i}$, where $h$ is 0 in the simulator and arbitrary but independent of the $L_{i}$s, $r$s, and $s_{i}$s in the true client (and $r$, $s_{i}$ are resampled for every such vector).
The following simple lemma shows that both such distributions are computationally indistinguishable from noise (and therefore from each other) conditioned on the LPN assumption, thus completing the proof.
@@
For $i\in[d]$, let $L_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{i}\times n_{i-1}}$. From this, for $j\in\{0,1,\dots,d\}$, let
$\mathcal{E}_{j}=\{L_{1}L_{2}\dots L_{j}r_{j}+\sum_{i=1}^{j}(L_{1}L_{2}\dots L_{i-1})s_{i}|r_{j}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{j}},s_{i}\leftarrow\mathcal{D}^{n_{i-1}}_{\mu_{i}}\}.$
Say that, for all $j\in[i]$, the distributions
$\mathcal{R}_{i}=\left\{r_{i-1}\overset{{\scriptscriptstyle\$}}{\leftarrow}R_{n_{j-1}}\right\},\mathcal{L}_{i}=\left\{L_{i}r_{i}+s_{i}:r_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}R_{n_{i}},s_{i}\leftarrow\mathcal{D}^{n_{i-1}}\right\}$ (1)
are $(T,\epsilon_{j})$-indistinguishable conditioned on $L_{j}$ (as given by the LPN assumption). Then $\mathcal{E}_{0}$ and $\mathcal{E}_{d}$ are $(T-\mathrm{\,poly}(n,\mu),\sum_{i=1}^{d}\epsilon_{i})$-indistinguishable conditioned on $(L_{1},\dots,L_{j})$. [^6] Note that the offset term represents the cost of the conversion step in the distinguisher.
@@
We prove the contrapositive via hybrid argument. Let there be an algorithm $\mathcal{A}$ distinguishing $\mathcal{E}_{0}$ and $\mathcal{E}_{d}$ with advantage $\geq\sum_{i=1}^{d}\epsilon_{i}$. Then there exists $j\in[d]$ s.t. $\mathcal{E}_{j-1}$ and $\mathcal{E}_{j}$ are distinguished by $\mathcal{A}$ with advantage $\geq\epsilon_{j}$. This implies an algorithm distinguishing the distributions $\mathcal{R}_{j}$ and $\mathcal{L}_{j}$ with advantage $\geq\epsilon_{j}$: sample matrices $L_{1},L_{2},\dots,L_{j-1}$ and noises $s_{1},\dots,s_{j-1}$ and apply them in the natural way to generate distributions $\mathcal{E}_{j-1}$ and $\mathcal{E}_{j}$, and then apply $\mathcal{A}$.
For the final security guarantee, we observe that the only place that the server can deviate from protocol without high-probability detection is in the last two queries (as all others are wrapped by Mat-Mul-Check. To check for $\alpha$-fraction of deviations in these queries, the client may introduce $\frac{c}{\alpha}$ multiplication queries with $B_{i}=0$. By linearity of the algorithm, the error introduced in either query propagates linearly into the result, and so if the server behaves dishonestly, the client will compute an incorrect result. The client may check this – $AB_{i}=0$ if $B_{i}=0$ – and given deviation rate $\alpha$, the probability of a deviation in $\frac{c}{\alpha}$ queries is $\geq 1-e^{-c}$, so any significant error rate can be detected cheaply and with very high probability.
@@ 7.4 Practical Improvements
Careful readers may note that this algorithm’s pseudorandom generation is very similar to a non-recursive instance of dual-LPN on the block syndrome matrix
@@
It may initially appear that we can simply substitute dual-LPN for this entire process and gain major client performance and simplicity improvements. We can, but only at factor-two (over unencrypted computation) cost to the server: the ancillary product of line 27 becomes much larger. We can, however, achieve nearly the same improvements by substituting dual-LPN for the final recursive step (i.e. replacing the generation $L_{d}v+s_{d}$ by $Ps_{d}$). This entirely eliminates the client dense multiplication, which is quite costly in regimes where $n$ is small. This does increase the server overhead, but this increase is on the multiplicative order of $\delta^{d-1}$. As a result, client performance improves substantially (and server performance decreases marginally) when $n$ is close to $n_{d}$, which occurs when $n$ is low or particularly high security is desired.
@@ 8 Experiments
We implemented Algorithm 2 in C++ and benchmarked it for both matrix-matrix multiplication and repeated matrix-vector multiplication. All experiments were run on a single core of an AMD Ryzen 5700X3D processor. All parameters were chosen for 128-bit security, using updated versions of the estimates of [LWYY24]. Results are in Tables 1 and 2. Client R. and Total R. indicate the ratios of client and total compute time, respectively, to the unencrypted compute time (denoted Local). We noticed substantially worse performance when $n$ was set to a power of two (presumably due to cache bottlenecking); therefore, we set $n$ to be one more than a power of two. Benchmarks with $n$ exactly a power of two are available in Tables 3 and 4 of the appendix; global performance is substantially worse, but relative performance is consistent.
$n$ Local (s) Client Pre. (s) Server Pre. (s) Client (s) Server (s) Client R. Total R.
@@
Figure 1: Matrix-Vectors performance benchmarks. Client and server times separated into amortized initialization and per-online-step. Average of five trials. Preprocessing times amortized across $n$ operations.
@@
Figure 2: Matrix-Matrix performance benchmarks. Average of five trials.
We see substantial performance improvements as $n$ grows in both contexts, but especially in matrix-vectors. Here, we achieve client improvements even at small $n$, with improvements increasing to factor-$25$ speedup for the client as we grow the matrix. Initialization costs are consistently negligible (due to the advantage of matrix-matrix over repeated matrix-vectors). In matrix-matrix, client improvements are less substantial, in large part due to the locality efficiencies available in dense (but not sparse) matrix multiplication. However, clients still achieve substantial advantage for medium-size matrices, up to nearly factor-three for $n=2^{14}$.
To the best of our knowledge, our method is the first to provide practical efficiency gains for the client in this parameter regime when encryption time is taken into account. Our efficiency is over two orders of magnitude higher than that of [BCH+24], despite encrypting both multiplicands while their approach encrypts one (note also that almost all of their cost is incurred by encryption/decryption). [^7] Their work, in turn, significantly outperforms [LZ22]. Our algorithm may be straightforwardly modified to only encrypt one multiplicand; this results in a significant additional improvement in client performance and server overhead.
Performance comparison to the later work of [BCH+25] is more complex. [^8] Our initial draft did not contain experimental data at their time of publication, so their work was unable to include experimental comparisons. We work over $\mathbb{Z}/2^{32}\mathbb{Z}$, while [BCH+25] work over $\mathbb{F}_{p}$ for $p$ a $32$-bit prime (and compares to plaintext operations over the same), rendering relative-to-plaintext performance not comparable. In addition, cryptographic assumptions need to be carefully evaluated to render accurate comparisons. Superficially, [BCH+25] obtain significant performance improvements for smaller values of $n$ ($\leq 2048$), while our work produces major gains for large $n$ ($\geq 4096$). In addition, [BCH+25] offers significant savings on client-side storage. The existence of constructions with tradeoffs suggests potential for future improvement.
@@ 9 Conclusion and Further Questions
We present natural definitions of Trapdoored-Matrix and Targeted-Trapdoored-Matrix objects, a method to turn instances of these objects into protocols for secure remote linear algebra, and instances of these objects that can be optimized to yield highly efficient versions of these protocols. It is not clear that these are the only (or best) primitives for this purpose, especially in the case of Trapdoored-Matrix, where evidence suggests the existence of notably faster pseudorandom matrices. Variants of our schemes may also be extensible to more general problems; harder problems in linear algebra, such as matrix polynomial evaluation, are a reasonable next step. Achieving a variety of operations parallel to that of [BCG+20a, CD23], or even [BCH+24] (while achieving practical efficiency gains) is a significantly more challenging problem, but would yield even broader applications.
@@ 9.0.1 Acknowledgements
Thanks to Alex Lombardi for helpful discussion on LPN and related topics.
Appendix 0.A Performance Analysis of Algorithm 2
@@ 0.A.1 Asymptotic Analysis
Fix $\delta,\epsilon,\lambda$. Let $\nu$ be as stated in Section 3. We assume $m,n\geq\nu$. Set $n_{i}=\delta n_{i-1}$ for some $\delta<\frac{1}{2}$, $\mu_{i}=n_{i-1}^{\epsilon-1}$, $n_{d}=\nu$. We have $\sum_{i=1}^{d}n_{i}\leq\frac{\delta n}{1-\delta}$, $\mu_{i}n_{i-1}\leq n_{i-1}^{\epsilon}$. We set $\lambda^{\prime}=\frac{\lambda}{\lg|R|}$ to achieve statistical $\lambda$-bit security in our Mat-Mul-Check calls.
To clarify the following expressions, we omit the costs of the many calls to Mat-Mul-Check in the intermediate expressions, as they are never dominant.
The preprocessing runtime is dominated by the computation of the products $L_{1}\dots L_{i}$, the Right-Hidden-Mult call (which dominates the init call), and the construction of $A^{\prime}$. The cost of the first is none (apart from the above check cost) to the client and $O(n^{\omega}\delta^{2\omega-3})$ to the server. The cost of the second is $O(\nu^{\omega-2}(m+n)(\delta n)+(m+n)n^{1+\epsilon})$ to the client and $O\left(n^{\omega}\delta^{\omega-1}+\min(m,\delta n)^{\omega-2}n\max(m,\delta n)\right)$ to the server. The construction of $A$ costs $O(mn\nu+mn^{1+\epsilon})$ to the client. [^9] In many expressions, $\nu$ can often be tightened to $\nu^{\omega-2}$ via fast matrix multiplication, but this is uninteresting and complicates the resulting expressions. In total, the client cost is $O\left(n\left(n^{\epsilon}(m+n)+\nu(\delta n+m)\right)\right)$ and the server cost is $O\left(n^{\omega}\delta^{2\omega-3}+\min(m,\delta n)^{\omega-2}n\max(m,\delta n)\right)$.
The online step runtime for the client is $O((m+n)l\nu+(m+n)n^{\epsilon}l)$ to construct $B^{\prime}$, $AB^{\prime}$, and $A^{\prime}(B+B^{\prime})$, and $O(ml)$ for the final subtractions. The server performs one $m\times n$-by-$n\times l$ multiplication in the final line and incurs additional cost $O(n\max(\delta n,l)\min(\delta n,l)^{\omega-2})$.

### v3

@@ 7.2 Security and Correctness Guarantees
Our protocol makes three guarantees. First, in the semi-honest model, the protocol will return exactly $AB$. Second, the protocol is computational zero-knowledge against even a dishonest server. Third, we obtain a weaker form of correctness of outputs: a server that behaves dishonestly in any setup query or fraction $\geq\alpha$ of the online invocations can be detected with high probability with $O\left(\frac{1}{\alpha}\right)$ additional queries (and if $l$ and $|R|$ are large, much more efficient per-call checks are available via Freivalds’ algorithm).
The first guarantee follows from expanding the return line and noting that it is equal to $AB$. We emphasize that all LPN noise exactly cancels; unlike some LWE-based protocols, we introduce no error in the output.
For the security guarantees, we first notice that Freivalds’ multiplication check (and therefore Protocol 7’s iterated version) will return true with probability 1 if the products are correct and probability $|R|^{-\lambda^{\prime}}$ otherwise, where $\lambda^{\prime}$ is the number of trials performed. Note that this provides low soundness error at very low computational cost for large rings: for 32-bit integers, for instance, we achieve soundness error $2^{-128}$ when $\lambda^{\prime}=4$. Assuming that $\lambda^{\prime}$ is picked so that this probability is negligible, we demonstrate that Protocol 7 is CZK.
@@
Given the LPN security assumption of Section 5, Protocol 7 is computational zero-knowledge for the server.
@@
We construct a server-view simulator. The simulator is quite simple: it runs a copy of the client with inputs $A=0,B_{i}=0$. The $L_{i}$ generated by the simulator are identically distributed to those generated by the true client. If the server misreports any of the products $L_{j}^{\top}L_{j-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i}$ or $A_{\mathbf{enc}}L_{1}L_{2}\dots L_{i}$, both the simulator and a true client will detect the error and terminate the protocol with probability $\geq 1-|R|^{-\lambda^{\prime}}$, so the complete views in this case are statistically indistinguishable.
Otherwise, letting $G^{(j)},S_{i}^{(j)}$ be the particular $G,S_{i}$ used in generation of $B^{\prime}$ in the $j$th online step, the rest of the client’s messages to the server can be concatenated into the matrix
$D+\langle L_{1}\dots L_{d}\rangle X+\sum_{i=0}^{d-1}\langle L_{1}\dots L_{i}\rangle Y_{i+1}$
where
$D$ $=\begin{bmatrix}A^{\top}&B^{(1)}&B^{(2)}&\dots&B^{(t)}\end{bmatrix}$
$X$ $=\begin{bmatrix}H^{\top}&G^{(1)}&G^{(2)}&\dots&G^{(t)}\end{bmatrix}$
$Y_{i}$ $=\begin{bmatrix}S_{i}^{\top}&T^{(1)}_{i}&T^{(2)}_{i}&\dots&T^{(t)}_{i}\end{bmatrix}$
Note in particular that $X\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{d}\times\mathrm{\,poly}(n)},Y_{i}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathcal{D}_{\mu_{i}}^{n_{i-1}\times\mathrm{\,poly}(n)}$. Then the server view is contained in the tuple
$\left((L_{i})_{i\in[d]},D+\langle L_{1}\dots L_{d}\rangle X+\sum_{i=0}^{d-1}\langle L_{1}\dots L_{i}\rangle Y_{i+1}\right)$
It therefore suffices to show that, for any $D,D^{\prime}$, the distribution (over $L_{i},X,Y$) of tuples of the form above is computationally indistinguishable from that of those of the form
$\left((L_{i})_{i\in[d]},D^{\prime}+\langle L_{1}\dots L_{d}\rangle X+\sum_{i=0}^{d-1}\langle L_{1}\dots L_{i}\rangle Y_{i+1}\right)$
(i.e. where we replace $D$ with $D^{\prime}$) for any $D,D^{\prime}$. The following lemma shows that any such distribution is indistinguishable from the distribution resulting from replacing the second tuple element with independent and uniform noise, and thus from that generated by any other $D^{\prime}$, completing the proof.
@@
Fix any $D$. For $j\in\{1,\dots,d\}$, let
$\mathcal{E}_{j}$ $=\left\{\left(\left(L_{i}\right)_{i\in[d]},D+L_{1}L_{2}\dots L_{j}X_{j}+\sum_{i=0}^{j-1}(L_{1}L_{2}\dots L_{i})Y_{i+1}\right)\middle|\right$.
$\left.\phantom{\Bigg(}L_{i}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{i-1}\times n_{i}},X_{j}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{j}\times\mathrm{\,poly}(n)},Y_{i}\leftarrow\mathcal{D}^{n_{i-1}\times\mathrm{\,poly}(n)}_{\mu_{i}}\right\}$
Let
$\mathcal{E}_{0}$ $=\left\{\left(\left(L_{i}\right)_{i\in[d]},D+U\right)\middle|L_{i}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{i-1}\times n_{i}},U\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{0}\times\mathrm{\,poly}(n)}\right\}$
$=\left\{\left(\left(L_{i}\right)_{i\in[d]},U\right)\middle|L_{i}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{i-1}\times n_{i}},U\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{0}\times\mathrm{\,poly}(n)}\right\}$
Under the assumption of Section 5 and assuming $n_{i}\geq\nu(\delta,\epsilon,\lambda),n_{i}\geq\delta n_{i-1},\mu_{i}\geq n_{i-1}^{\epsilon-1}\forall i\in[d]$, $\lambda=\Omega(\lg n)$, $\mathcal{E}_{d}$ is $(\lambda-\lg d)$-bit computationally indistinguishable from the noise distribution $\mathcal{E}_{0}$.
@@
We prove the contrapositive. Let there be an efficient algorithm $\mathcal{A}$ distinguishing $\mathcal{E}_{0}$ and $\mathcal{E}_{d}$ with advantage $\geq\sum_{i=1}^{d}\epsilon_{i}$. Then there exists $j\in[d]$ s.t. $\mathcal{E}_{j-1}$ and $\mathcal{E}_{j}$ are distinguished by $\mathcal{A}$ with advantage $\geq\epsilon_{j}$. Then there exists an efficient algorithm distinguishing the distributions
$\left\{(L_{j},L_{j}X_{j}+Y_{j})\middle|L_{j}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{j-1}\times n_{j}},X_{j}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{j}\times\mathrm{\,poly}(n)},Y_{j}\leftarrow\mathcal{D}^{n_{j-1}\times\mathrm{\,poly}(n)}_{\mu_{j}}\right\}$
and
$\left\{(L_{j},U)\middle|L_{j}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{j-1}\times n_{j}},U\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{j-1}\times\mathrm{\,poly}(n)}\right\}$
with advantage $\epsilon_{j}$: enhance samples from the former into samples from $\mathcal{E}_{j}$ and samples from the latter into samples from $\mathcal{E}_{j-1}$ by drawing the other random variables and computing the second tuple element from the results, and then apply $\mathcal{A}$. This shows that the LPN assumption does not hold.
For the final security guarantee, we observe that the only place that the server can deviate from protocol without high-probability detection is in line 4 of the online multiplication subroutine (as all other requests are checked by Protocol 7). To check for an $\alpha$-fraction of deviations in these queries, the client may introduce $\frac{c}{\alpha}$ multiplication queries with $B_{i}=0$. By linearity of the downstream sections of the protocol, the error introduced in either query propagates linearly into the result through $H$ and the $S_{i}$, and so if the server behaves dishonestly, the client will compute an incorrect result with high probability. The client may check this – $AB_{i}=0$ if $B_{i}=0$ – and given deviation rate $\alpha$, the probability of a deviation in at least one of the $\frac{c}{\alpha}$ checking queries is $\geq 1-e^{-c}$, so any significant error rate can be detected cheaply and with high probability.
@@ 7.4 Practical Improvements
Careful readers may note that this protocol’s pseudorandom generation is very similar to a non-recursive instance of dual-LPN on the block syndrome matrix
@@
It may initially appear that we can simply substitute dual-LPN for this entire process and gain major client performance and simplicity improvements. We can, but only at constant-factor (over unencrypted computation) cost to the server, as the server would need to multiply the syndrome matrix by $A_{\mathrm{enc}},B_{\mathrm{enc}}$. We can, however, achieve nearly the same improvements by substituting dual-LPN for the final recursive step (i.e. replacing the generation $L_{d}G+T_{d}$ by $PT^{\prime}_{d}$). This entirely eliminates the client dense multiplication, which is quite costly in regimes where $n$ is small. This does increase the server overhead, but this increase is on the multiplicative order of $\delta^{d-1}$. As a result, client performance improves substantially (and server performance decreases marginally) when $n$ is close to $n_{d}$, which occurs when $n$ is low or particularly high security is desired.
@@ 8 Experiments
We implemented Protocol 7 in C++ and benchmarked it for both matrix-matrix multiplication and repeated matrix-vector multiplication. Experiments ran on a single core of an AMD Ryzen 5800X3D processor. Parameters were chosen for 128-bit security, using updated versions of the estimates of [29]. Results are in Tables 1 and 2. Client R. and Total R. denote the ratios of client and total compute time, respectively, to the unencrypted compute time (denoted Local). We noticed substantially worse performance when $n$ was a power of two, presumably due to cache associativity; therefore, we set $n$ to be one more than a power of two. Benchmarks with $n$ exactly a power of two are available in Tables 3 and 4 of the appendix; global performance is substantially worse, but relative performance is fairly consistent.
$n$ Local (s) Client Init. (s) Server Init. (s) Client (s) Server (s) Client R. Total R.
@@
Table 1: Matrix-Vectors performance benchmarks. Client and server times separated into amortized initialization and per-online-step. Average of five trials. Initialization times amortized across $n$ operations. Note that some server time is conservatively attributed to the client in simulation, so server times are slightly below naive time for small operand sizes.
@@
Table 2: Matrix-Matrix performance benchmarks. Average of five trials.
We see substantial performance improvements as $n$ grows in both contexts, but especially in matrix-vectors. Here, we achieve client improvements even at small $n$, with improvements increasing to factor-$25$ speedup for the client as we grow the matrix. Initialization costs are consistently negligible (due to the advantage of matrix-matrix over repeated matrix-vectors). In matrix-matrix, client improvements are less substantial, in large part due to the locality efficiencies available in dense (but not sparse) matrix multiplication. However, clients still achieve substantial advantage for medium-size matrices, up to nearly factor-three improvement for $n=2^{14}$.
To the best of our knowledge, our method is the first to provide practical efficiency gains for the client in this parameter regime when encryption time is taken into account. Our efficiency is over two orders of magnitude higher than that of [5], despite encrypting both multiplicands while their approach encrypts one (note also that almost all of their cost is incurred by encryption/decryption). [^7] Their work, in turn, significantly outperforms [30]. Our protocol may be straightforwardly modified to only encrypt one multiplicand; this results in a significant additional improvement in client performance and server overhead.
Performance comparison to the later work of [7] is more complex. [^8] Our initial draft did not contain experimental data at their time of publication, so their work was unable to include experimental comparisons. We work over $\mathbb{Z}/2^{32}\mathbb{Z}$, while [7] works over $\mathbb{F}_{p}$ for $p$ a $32$-bit prime (and compares to plaintext operations over that field), rendering relative-to-plaintext performance incomparable. In addition, results are highly sensitive to deviations in exact security assumptions. With these caveats, [7] obtains significant performance improvements for smaller values of $n$ ($\leq 2048$), consistent with the increasing importance of some of their overhead improvements at those scales, while our methods are more efficient for larger $n$. Their methods also offer significant savings on client-side storage for matrix-vectors. The existence of constructions with tradeoffs suggests potential for future improvement.
@@ 9 Conclusion and Further Questions
We present natural definitions of Trapdoored-Matrix and Targeted-Trapdoored-Matrix objects, a method to turn instances of these objects into protocols for secure delegated linear algebra, and optimized constructions of all of the above. More recently, other constructions of the above primitives have been put forward, and the existence of still better constructions remains likely. Variants of our schemes may also be extensible to more general problems; harder problems in linear algebra, such as matrix polynomial evaluation, are a reasonable next step. Achieving a variety of operations comparable to those of [11, 14, 5] (while maintaining efficiency gains) seems significantly more challenging, but would yield even broader applications.
@@ 9.0.1 Acknowledgements
Thanks to Alex Lombardi for helpful discussion on LPN and related topics. Thanks to the TCC’25 reviewers for helpful comments and suggestions.
Appendix 0.A Performance Analysis of Protocol 7
@@ 0.A.1 Asymptotic Analysis
Fix $\delta<\frac{1}{2},\epsilon,\lambda$. Let $\nu$ be as stated in Section 3. We assume $m,n\geq\nu$. Set $n_{i}=\delta n_{i-1}$, $\mu_{i}=n_{i-1}^{\epsilon-1}$, $n_{d}=\nu$. We have $\sum_{i=1}^{d}n_{i}\leq\frac{\delta n}{1-\delta}$, $\mu_{i}n_{i-1}\leq n_{i-1}^{\epsilon}$. We set $\lambda^{\prime}=\frac{\lambda}{\lg|R|}$ to achieve statistical $\lambda$-bit security in multiplication checks.
For simplicity, we omit the costs of the many calls to Protocol 7, as they are never dominant in reasonable parameter regimes. They are included in Appendix 0.A.2, but it suffices to observe that they are dominated by $O(1)$ multiplications of $O(m+n)\times O(n)$ matrices by $O(n)\times O(\lambda^{\prime})$ matrices.
Initialization:
The cost to compute all product sequences of the $L_{i}$s is $O(n^{\omega}\delta^{\omega-1})$ to the server (as this step is dominated up to constants by a $\delta n\times n$-by-$n\times\delta n$ multiplication, with cost $\frac{1}{\delta}(\delta n)^{\omega}$). The cost to compute $A^{\prime}$ is $O(mn\nu^{\omega-2})$ for the first operation and $O(mn\sum_{i=0}^{d}(n\delta^{i})^{\epsilon})=O(mn^{1+\epsilon})$ for the second. The cost to compute the partial products with $A_{\mathbf{enc}}$ is $O\left(\min(m,\delta n)^{\omega-2}n\max(m,\delta n)\right)$. The cost to compute the partial products of $A^{\prime}$ with the various $L_{i}$ is dominated by the cost to compute $A^{\prime}$ itself. Total cost (omitting check terms) is therefore $O(n^{\omega}\delta^{\omega-1}+\min(m,\delta n)^{\omega-2}n\max(m,\delta n))$ to the server and $O(n(m+n)(n^{\epsilon}+\nu^{\omega-2}))$ to the client.
Multiplication:
Computation of $B^{\prime},AB^{\prime}$ costs $O(l(m+n)(\nu+n^{\epsilon}))$ to the client (we do not get $\nu^{\omega-2}$ as $l$ may be smaller than $\nu$; the bounds can be improved by splitting on this event). Computation of $A_{\mathbf{enc}}B_{\mathbf{enc}}$ costs $O(mnl\min(m,n,l)^{\omega-3})$, and partial products cost $O(n\max(l,\delta n)\min(l,\delta_{n})^{\omega-2})$ to the server. Computation of $A^{\prime}B_{\mathbf{enc}}$ costs $O(lm(\nu+n^{\epsilon}))$ to the client. Then the total cost is, apart from the large multiplication, $O(n\max(l,\delta n)\min(l,\delta n)^{\omega-2})$ to the server and $O(l(m+n)(\nu+n^{\epsilon}))$ to the client.